The admission fee at an amusement park is $3.50 for children and $6.00 for adults. On a certain day, 341 people entered the park, and the admission fees collected totaled 1606 dollars. How many adults were admitted?
Number of children equals
Number of adults equals

Answer by Cromlix(4381)

Hi there,
Number of children = x
Number of adults = y
x + y = 341......(1)
3.50x + 6.00y = 1606.......(2)
Multiply Eq(1) by 6
6x + 6y = 2046.....(1)
3.5x + 6.00y = 1606.....(2)
Subtract Eq(2) from Eq(1)
2.5x = 440
x = 176
Substitute x = 176 in Eq(1)
x + y = 341.....(1)
176 + y = 341
y = 341 - 176
y = 165.
Number of children admitted = 176
Number of adults admitted = 165
Hope this helps :-)